Fermi National Laboratory


Accelerator Update for December 2, 2005 - December 5, 2005

Friday December 2

The day shift began with the Tevatron (TeV) sector B17 work continuing, and with MiniBooNE, NuMI, and the Meson MTest and MCenter experiments taking beam.

During the day shift, cryo system experts replaced a bad valve, a Main Injector (MI) expert began a four hour 2.5 Mhz study, Antiproton Source (Pbar) experts conducted reverse proton studies, and a Booster power supply (GMPS) tripped off; it reset okay.

On the evening shift Pbar continued their reverse protons studies and MI conducted NuMI intensity/stability studies.

Saturday December 3

The midnight shift began with TeV B17 work continuing, with Pbar conducting reverse proton studies, and with MiniBooNE, NuMI, MTest, and MCenter taking beam.

Pbar ended their studies at 7:30 AM. Operators discovered a Booster LCW leak around 8 AM. Water technicians had the leak repaired by noon, but a 400 MeV area Debuncher problem held beam off for another hour. Pbar and NuMI conducted studies until two Booster RF stations (BRF15 & 16) began sparking due to a vacuum problem. Vacuum techs accessed Booster and found a vacuum leak on BRF16. They repaired the leak and beam had resumed by midnight.

Sunday December 4

The midnight shift began with TeV B17 work continuing, with Pbar conducting reverse proton studies, and with MiniBooNE, NuMI, MTest, and MCenter taking beam.

Pbar studies ended at 3:30 AM. TeV experts began conducting a B1 hipot around 8 AM.

During the day and evening shifts, Pbar experts conducted AP2 transfer line studies and MI resumed their 2.5 Mhz studies.

Monday December 5

The midnight shift began with TeV B17 work continuing, with Pbar studies, and with MiniBooNE, NuMI, MTest, and MCenter taking beam.

A NuMI power supply (V118) tripped off. Power supply experts replaced the supply. By the end of the shift, experts began the mechanical and cryo system walkthrough of TeV sector B1.

The Plans for Monday

The plans are to continue with the TeV B1 recovery. The mechanical and cryo system signoff should occur around noon today, with the cooldown beginning soon after. The separator bake out should be finished on Tuesday and cool by Wednesday, 12/7/05.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 12/5/2005   email Fermilab